Suppose you are considering purchasing a $7,500 machine that will increase sales by $2,000 per year over a 4-year period. At the end of the 4 year period, the machine will have a salvage value of 0. The asset will be depreciated to 0 using straight line method. If the appropriate tax rate of 32%, and a required return of 9%, what is the NPV of the project?
Answer #1

Annual depreciation = 7,500 / 4 = 1,875

Operating cash flow = (Revenue - depreciation)(1 - tax) + depreciation

Operating cash flow = (2,000 - 1,875)(1 - 0.32) + 1,875

Operating cash flow = 85 + 1,875

Operating cash flow = 1,960

NPV = Present value of cash inflows - present value of cash outflows

NPV = -7500 + 1,960 / (1 + 0.09)1 + 1,960 / (1 + 0.09)2 + 1,960 / (1 + 0.09)3 + 1,960 / (1 + 0.09)4

NPV = $1,150.15